A clown in the park gave a child a balloon with 6.00 L of helium at 280 K. What will the volume of the balloon be when the child takes it into a heated house at 22ºC?



1
Expert's answer
2021-04-12T03:45:38-0400

V1/T1 =V2/T2

= 6.00/280 =X/295

= 0.0214 = X/295

X = 295 × 0.0214

= 6.3214L